OpenStudy (anonymous):

Identify the x and y intercepts for the equation 4x – 3y = -12.

OpenStudy (anonymous):

A(4, 0) and (0, –3) B(3, 0) and (0, –4) C(4, 0) and (0, 3) D(–3, 0) and (0, 4) E(3, 0) and (0, 4)

OpenStudy (anonymous):

First, put it into slope-intercept form

OpenStudy (radar):

Here is how you work them. First let x = 0 and solve for y that gives you (0,y) Next let y = 0 and solve for x, that gives you (x,0) That is the intercepts.

OpenStudy (radar):

x=0, 0-3y=-12, so y=4 (0,4) y=0, 4x-0=-12, 4x=-12, x=-3, (-3,0)

OpenStudy (radar):

Do you see which one it is ? I see it.

OpenStudy (anonymous):

i just solved but thnks lol u gave me the answer

OpenStudy (radar):

It is pretty easy that way don't you think?

OpenStudy (anonymous):

yea a lot easier

OpenStudy (radar):

Good luck with your studies.

OpenStudy (anonymous):

thnks